21. Two of your friends, Rex and Cody, are having a disagreement. Rex argues that 0.020 M HCl is more acidic than 5.0 M CH3CH2COOH (pKa = 4.87) because nitric acid is a strong acid. Cody argues, however, that 5.0 M CH3CH2COOH (pKa = 4.87) is more acidic than 0.020 M HNO3 because it has a higher concentration. Who is correct? Prove your point by calculating both pHs.


1
Expert's answer
2021-08-02T12:26:18-0400

Solution

"CH_3CH_2COOH(pKa=4.87)"


pKa = 4.87


logka= -4.87


ka = "1.34\times10^{-5}"


"[H^+]=\sqrt{1.34\times10^{-5}\times5.0}"


"=8.19\times10^{-3}"


pH= -log(8.19"\times10^{-3})" = 2.09


pH of HCl"=log(0.020)=1.70"


pH of HN03"=log(0.020)=1.70"


Rex is correct propanoic acid is weaker compared to nitric acid because the lower the value of pH the stronger the acid.